Logiweb bibliography of ijcar base

Up Help

[0] ijcar base

The pyk compiler, version 0.grue.20060417 by Klaus Grue GRD-2006-02-24.UTC:10:23:46.350024 = MJD-53790.TAI:10:24:19.350024 = LGT-4647493459350024e-6